OpenStudy (anonymous):

A box contains only red yellow and green jelly beans. The number of red jelly beans 4/5 of the number of green jelly beans 3/4 the number of yellow ones. If there are 470 beans in total, what's the probability of picking yellow?

OpenStudy (anonymous):

I think I know this: what is 470 divided by 4?

OpenStudy (anonymous):

How is that

OpenStudy (anonymous):

470/4=235/2=117.5 @julianassange